Shift report 20010922 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Sep 22 2001 - 09:03:47 EDT

  • Next message: Apache: "Shift report 20010922 08:00-16:00"

    ------------------------------------------------------------
    Date: Sat Sep 22 09:03:47 2001, Shift: 00:00-08:00
    Supervisor: hito, with: 
    Log book#9 pages 24 - 25
    
    Shift Summary:
    ------------------------------------------------------------
    1.  beam until 0:20  
        At 0:20 QLI.
        At 1:34, power dip.  
    2.  None
    3.  Because of the power dip, we completly lost the
    communication to all 4032.  I(hiro) went into the IR
    in the control access to check the HV.  It was still
    on.  Until it is fixed, I suggests that nobody will 
    recyle the 4032s (BBR, BBL, BBBig, Tile, gating grid,
    , tofwcal).
    T3 first module still needs fix.
    Magnet is in standy-fault mode, the magnet watch needs 
    to fix it.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Sep 22 2001 - 09:04:22 EDT